OpenStudy (anonymous):

best answer receives medal and fan :):):):):O Michelle and Maggie are at baseball practice. Michelle throws a ball into the air and when it drops to a height of 5ft., she hits the ball. The height of the ball is modeled by the graph below where t = time in seconds and h = height of the ball from the ground. Maggie is throwing a ball into the air and catching it. The height of Maggie’s ball is modeled by the function h(t) = –16t2 + 48t + 15.

OpenStudy (anonymous):

Maggie is throwing a ball into the air and catching it. The height of Maggie’s ball is modeled by the function h(t) = –16t2 + 48t + 15. Part 1. Which ball goes higher in the air, the ball that is hit or the ball that is thrown? Use complete sentences and show all work to explain how you determined the height that each ball reaches. Part 2. Determine which girl is likely to be standing on a raised platform. Use complete sentences to explain how you determine which girl is on the platform and then determine the height of the platform. Part 3. Which ball is traveling at a faster average rate of change on the way up? Use complete sentences to explain how you determined the interval at which the height of the ball is increasing and the average rate of change.

OpenStudy (anonymous):

where is the graph?

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

@dumbsearch2

OpenStudy (anonymous):

@poopsiedoodle

OpenStudy (anonymous):

for Maggie h(t)=-16t^2+48t+15\[=-16(t^2-3t+\frac{ 9 }{4 }-\frac{ 9 }{4 })+15=-16\left( t-\frac{ 3 }{ 2 } \right)^2+16*\frac{ 9 }{4 }+15\] \[=-16\left( t-\frac{ 3 }{2 } \right)^2+51\] maximum height is 51 and is attained at t=3/2

OpenStudy (anonymous):

Mitchel's ball reaches to a height of 22.5 only

OpenStudy (anonymous):

okay now what about part b?

OpenStudy (anonymous):

b. for Maggie put t=0 and find h(0) for Mitchel is standing on the ground and hits the ball at height of 5

OpenStudy (anonymous):

have you calculated h(0),just plug t=0

OpenStudy (anonymous):

\[h(0)=-16(0)^2+48*0+15=?\]

OpenStudy (anonymous):

where are you?

OpenStudy (anonymous):

im trying to calculate h(0)=

OpenStudy (anonymous):

im getting a lil confussed though

OpenStudy (anonymous):

it is 15 so Maggie is on a raised platform and the height of platform is 15

OpenStudy (anonymous):

ooo okay I see

OpenStudy (anonymous):

part 3?

OpenStudy (anonymous):

c.for Mitchel when t=o=0, h=5 when t=1,h=22.5 \[Rate~of~change=\frac{ 22.5-5 }{1-0 }=17.5ft/\sec.\] for Maggie,\[when~t=0,h=15~ft,when~t=\frac{ 3 }{ 2 }\sec.,h=51 ft\] \[Average~Rate~of~change=\frac{ 51-15 }{\frac{ 3 }{ 2 }-0 }=36*\frac{ 2 }{3 }=24ft/\sec\] so Maggie's ball is rising at a faster average rate.

OpenStudy (anonymous):

can you explain the steps?

OpenStudy (anonymous):

just note the height of ball from the graph when t=0 and t=1 difference is change in height. speed=(change in height)/time

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

yw